OpenStudy (anonymous):

Molarity question: you have 695 mL of a o.61 M solution of nitric acid. what is the weight of the nitric acid in solution. 27 g

OpenStudy (aaronq):

so you have 0.61 M M= n/L ... so you have 0.61 moles of HNO3 in 1 L of water.. 0.61 mol/L x 63g/mol = 38.43 g/L you have a volume of 0.695 L.. 38.43 g/L x 0.695 L = 26.71 g of HNO3

OpenStudy (anonymous):

what about 27g? I dont need to use this mass?

OpenStudy (aaronq):

i think that was the answer

OpenStudy (anonymous):

Thank you :) could you help with other question? 66.32 g Pb(So4)2 are dissolved into a 1350 mL solution and 136.2 g FeSo4 are dissloved into a 980 mL solution. which solution contains a higher molarity of sulfate ion?

OpenStudy (aaronq):

ill tell you how to do it, how bout that? find the number of moles then find molarity .. wait are you sure it's Pb(SO4)2? not PbSO4

OpenStudy (aaronq):

so yeah, find moles in each solution.. n=m/MW then find molarity M=n/L which ever one is higher wins

OpenStudy (anonymous):

thanks, MW stands for mass weight?

OpenStudy (aaronq):

molecular weight or molecular mass